OpenStudy (anonymous):

x + 6 - 5 = 8 - 2 x=?

OpenStudy (anonymous):

let me see

OpenStudy (anonymous):

5 is the answer

OpenStudy (anonymous):

x+6-5=8-2 x+1=6 -1 -1 _______________________________ x= 6-1=?

OpenStudy (anonymous):

cus 5+6-5=6

OpenStudy (anonymous):

^^^ You didnt have to move anything over but the -1